# Extended Friction Models for the Physics Simulation of Servo Actuators

### v2

@@ II-A Test bench dynamics
where $\theta$ is the servo actuator position, $\tau_{m}$ the motor torque, $\tau_{e}(\theta)=mgl\sin(\theta)$ the external torque caused by gravity, $\tau_{f}$ the torque caused by friction and $J$ the inertia. The inertia $J$ is the sum of the load inertia $ml^{2}$ and the servo actuator apparent inertia $J_{m}$ (sometimes referred to as armature). When the rotor inertia $J_{r}$ is known from constructor data, $J_{m}$ is given by $J_{m}=N^{2}J_{r}$ with $N$ the inverse of the gear ratio [17]. Otherwise, $J_{m}$ is an extra parameter to identify. In the following notations, we drop the dependency of $\tau_{e}$ to $\theta$ for clarity.
@@ II-B Friction model
Thus, the resistive torque $\tau_{f}$ caused by friction is given by $\tau_{f}^{stop}$, as computed in equation (4), and limited to the range $[-\tau_{f}^{m},\tau_{f}^{m}]$:
@@ IV-B Voltage control law
where $U_{max}$ is the motor maximum voltage and $clip$ a function that bounds a value to a given interval. The motor torque can then be derived from DC motor equations [17]:
@@ IV-D Simulating the pendulum test bench
3 $\tau_{e,k}\leftarrow mglsin(\theta_{k})$; test bench dynamics
@@ VI-B Validation on 2R arms
Each trajectory is recorded and simulated using both high and low proportional gains. High gains provide precise control, while low gains offer loose control, as typically used in RL. Fig. 1 details a log obtained with an MX-106 on a triangular wave with low gains.

### v4

@@ II-A Test bench dynamics
where $\theta$ is the servo actuator position, $\tau_{m}$ the motor torque, $\tau_{e}(\theta)=-mgl\sin(\theta)$ the external torque caused by gravity, $\tau_{f}$ the torque caused by friction and $J$ the inertia. The inertia $J$ is the sum of the load inertia $ml^{2}$ and the servo actuator apparent inertia $J_{m}$ (sometimes referred to as armature). When the rotor inertia $J_{r}$ is known from constructor data, $J_{m}$ is given by $J_{m}=N^{2}J_{r}$ with $N$ the inverse of the gear ratio [17]. Otherwise, $J_{m}$ is an extra parameter to identify. In the following notations, we drop the dependency of $\tau_{e}$ to $\theta$ for clarity.
@@ II-B Friction model
Thus, the resistive torque $\tau_{f}$ caused by friction is given by $\tau_{f}^{stop}$, as computed in equation (4), and limited to the range $\tau_{f}^{m},\tau_{f}^{m}]$:
@@ IV-B Voltage control law
where $U_{max}$ is the motor maximum voltage and clipclip a function that bounds a value to a given interval. The motor torque can then be derived from DC motor equations [17]:
@@ IV-D Simulating the pendulum test bench
3 $\tau_{e,k}\leftarrow-mglsin(\theta_{k})$; test bench dynamics
@@ VI-B Validation on 2R arms
Each trajectory is recorded and simulated using both high and low proportional gains. High gains provide precise control, while low gains offer loose control, as typically used in RL. Fig. 1 details a log obtained with the Dynamixel arm on a triangular wave with low gains.